OpenStudy (anonymous):

find the eighth term of a geometric sequence if the second term is 3 and the fourth term is 27,r>0

OpenStudy (anonymous):

ar = 3 ar^3 = 27

OpenStudy (anonymous):

r^2 = 27/3 = 9 r = 3 since r > 0

OpenStudy (anonymous):

a(3) = 3 a = 1 ar^7 = 1(3)^7

OpenStudy (anonymous):

2187

OpenStudy (cwrw238):

ar^3 / ar = r^2 = 27/3 = 9 so r = 3

OpenStudy (cwrw238):

8th term - as moneybird said